# Towards heterogeneous parallelism for SPHinXsys

### 12 Mar 2026 v1

@@ Abstract
This paper presents a Weakly Compressible Smoothed Particle Hydrodynamics (WCSPH) method solving the two-equation Reynolds-Averaged Navier-Stokes (RANS) model for the turbulent wall-bounded flows with or without flow separation. The inconsistency between the Lagrangian nature of the SPH and RANS model, primarily caused by intense shearing and near-wall discontinuities, is firstly revealed and addressed by the improved mainstream and near-wall treatments, respectively. The mainstream treatments, including Adaptive Riemann-eddy Dissipation (ARD) and de-noised transport velocity formulation, address dissipation incompatibility, turbulent kinetic energy disturbance and over-prediction issues. The near-wall treatments, such as the particle-based wall model realization, weighted near-wall compensation scheme, and constant wall-normal spacing strategy, improve the accuracy and stability of the adopted wall model, where the wall dummy particles are still used for future coupling of solid dynamics. Besides, to perform rigorous convergence tests, a level-set-based Boundary-Offset Technique (BOT) is developed to ensure consistent wall-normal distance across different resolutions. Several benchmark wall-bounded turbulent flow cases are simulated, including straight, mildly curved, strongly curved, Half Converging-Diverging (HCD) channels, and a fish-pass. The present method yields smoothed and reasonably accurate results, and, to the best of our knowledge, achieves for the first time satisfactory convergence of both velocity and turbulent kinetic energy in SPH-RANS simulations. The proposed method bridges particle-based and mesh-based RANS models, providing adaptability for other turbulence models and potential for turbulent fluid-structure interaction (FSI) simulations.
@@
Smoothed particle hydrodynamics , Turbulence , RANS , Wall-bounded flow , Lagrangian

### 17 Mar 2026 v2

@@ Abstract
Simulations based on particle methods, such as Smoothed Particle Hydrodynamics (SPH), are known to be computationally demanding. While such methods have for long been executed in parallel on multi-core CPUs, in recent years the increasing adoption of many-core accelerators, such as GPUs. However, hardware fragmentation and vendor-specific programming interfaces are still characterizing their market. Hence, support for various hardware configurations may easily lead to non-trivial and less maintainable implementations. To leverage over some higher-level specifications have become available recently, such as the SYCL programming standard, this work highlights the initial effort in adopting the SYCL standard for the execution of SPHinXsys, an open-source multi-physics library. The result is an execution model able to run the same implementation on variable (heterogeneous) hardware, with considerable speed-up compared to the current multi-core CPU parallelization. Among others, representation of data-structures for parallel access, communication strategies, and parallel methods for data sorting will be topics discussed in depth. Benchmarks has also been presented, showcasing performance comparisons between the current multi-core CPU implementation and the newly introduced SYCL parallelization with a GPU back-end.
@@
Smoothed particle hydrodynamics , GPU acceleration , SYCL , data-structure